OpenStudy (anonymous):

Evaluate the Indefinite Integral 2u(u^2+1)^2 du

OpenStudy (amistre64):

D(u^2 +1 ) = 2u

OpenStudy (anonymous):

Try using substitution and let x = u^2 + 1

OpenStudy (amistre64):

lol.......trig sub ;)

OpenStudy (anonymous):

not so much trig sub but regular substitution

OpenStudy (anonymous):

x = u^2 + 1 dx = 2u du \[\int\limits_{}^{}x^2dx = \frac{1}{3}x^3 = \frac{1}{3}(u^2 + 1)^3\]

OpenStudy (amistre64):

...trig sub!!... its all I can do ;)

OpenStudy (anonymous):

amistre loves his trig subs

OpenStudy (amistre64):

trig sub tis; trig sub that; but the trig sub tshirt and ride the rides.....trig sub!!!!

OpenStudy (anonymous):

ROFLLL!!

OpenStudy (anonymous):

chiaroscuro - was that your final answer? It's actually a multiple choice question and I can't seem to figure out if I got the right answer or if it's "none of the above"

OpenStudy (anonymous):

1/3 (u^2 + 1)^3